Bavarian Crystal Works designs and produces lead crystal wine decanters for export to international markets. The production manager of Bavarian Crystal Works estimates total and marginal production costs to be T C = 10, 000 + 40Q + 0.0025Q2 MC = 40 + 0.005Q where revenues are measured in U. S. dollars and Q is annual decanter production. a. What is the optimal level of production of wine decanters? What is the marginal revenue from the last wine decanter sold? b. What are the total revenue, total cost, and net benefit (profit) from selling the optimal number of wine decanters? c. Is the optimal profit in answer b good enough to keep Bavarian Crystal Works in the market? Please comment on it.
Answer #1

ANSWER:-

(a)

Optimal quantity is where

MR=MC

40 + 0.005Q=70

Q=6000

The correct answer is: 6000

What is the marginal revenue from the last wine decanter sold?

The correct answer is: $ 70

(b)

it’s a perfectly competitive market so, P=MC

TOTAL REVENUE=PQ

=6000*70

TR=$420,000

The correct answer is: $420,000

Answer:- TC=10,000 + 40Q + 0.0025Q2

10,000 + 40(6000) + 0.0025(6000)2

=$340,000

The correct answer is: $340,000

(c)

Net Benefit= Total revenue-total cost

=$420,000- $340,000

=$80,000

The correct answer is: $80,000